# Physical Degradation Model-Guided Interferometric Hyperspectral Reconstruction with Unfolding Transformer

### v1

@@ Abstract
Interferometric Hyperspectral Imaging (IHI) is a critical technique for large-scale remote sensing tasks due to its advantages in flux and spectral resolution. However, IHI is susceptible to complex errors arising from imaging steps, and its quality is limited by existing signal processing-based reconstruction algorithms. Two key challenges hinder performance enhancement: 1) the lack of training datasets. 2) the difficulty in eliminating IHI-specific degradation components through learning-based methods. To address these challenges, we propose a novel IHI reconstruction pipeline. First, based on imaging physics and radiometric calibration data, we establish a simplified yet accurate IHI degradation model and a parameter estimation method. This model enables the synthesis of realistic IHI training datasets from hyperspectral images (HSIs), bridging the gap between IHI reconstruction and deep learning. Second, we design the Interferometric Hyperspectral Reconstruction Unfolding Transformer (IHRUT), which achieves effective spectral correction and detail restoration through a stripe-pattern enhancement mechanism and a spatial-spectral transformer architecture. Experimental results demonstrate the superior performance and generalization capability of our method.
@@ 2.2 Degradation Modeling and Dataset Synthesis
Learning-based restoration approaches typically draw their potential from data. However, the acquisition of large-scale real datasets is often challenging, and one solution is using degradation models as priors to generate synthetic datasets. Common models includes Gaussian white noise and homoscedastic Gaussian noise [33, 1, 16], among others. Wei [36, 37] introduced a camera calibration-based, precise noise model that allows simulated data to yield low-light noise recovery akin to real-data training outcomes. Zhang[47, 48] extended this to HSI denoising, presenting a realistic dataset and an HSI noise model. These methods provide inspirations for addressing the challenges in existing IHI calibration and generating training data.
@@ 2.3 Hyperspectral Image Reconstruction
Existing HSI reconstruction focus on tasks such as RGB-to-HSI conversion and compressed sensing imaging (SCI) [41, 2]. Deep learning approaches outperforms traditional model-driven methods [5, 42] in terms of performance and generalization. These methods are categorized into two types: end-to-end (E2E) and model-aided networks. E2E methods directly learn the reconstruction mapping, with examples including CNN-based TSA-Net [29] and HDNet [19], RNN-based BIRNAT [11], and transformer-based MST, MST++ , and CST [7, 8, 6]. Model-aided networks, exemplified by Deep Plug-and-Play (PnP) [49] and Deep Unfolding (DU), integrate iterative optimization with deep learning. Among these, DU adopts the same training strategy as E2E methods, thereby enhancing the network’s adaptability to iterations compared to PnP, and achieving superior restoration performance. Examples include GAP-Net [30], transformer-based DAUHST [9], Proximal Gradient Descent (PGD)-based RDLUF [13], and Memory-Augmented MAUN [18]. However, for IHI tasks, these algorithms have yet to be applied and optimized.
@@ 3.1 Degradation Modeling
The input light signal can be considered as a hyperspectral signal in a certain wavelength range $[\lambda_{\text{min}},\lambda_{\text{max}}]$. Assuming expression of the signal in the form of wavenumber $\nu$ ($\nu=1/\lambda\in[1/\lambda_{\text{max}},1/\lambda_{\text{min}}]$) is $\mathbf{B}_{0}(\nu)\in\mathbb{R}^{H\times W\times N}$. According to theories of IHI [4], $\mathbf{B}_{0}$ is Fourier transform-related to the interferogram. However, due to the imaging system’s varying sensitivity to light in different frequencies, the actual captured spectral signal is the result of the original signal multiplied by a coefficient vector $\mathbf{A}$ in the spectral domain. Additionally, pose deviations in the interferometer lead to asymmetry in the interferogram, introducing a complex component as phase error into $\mathbf{A}$. We define $\mathbf{A}\in\mathbb{C}^{W\times N}$ as the absolute response coefficient, and the interferogram signal can be represented as:
@@ 4.1 Datasets and Settings
Synthetic Datasets. We prepare two HSI source datasets for simulation, namely HSOD-BIT [34] and Houston. HSOD-BIT is a multi-spectral HSI dataset consists of 319 fine-quality HSIs originating from objective detection. The dataset possess suitable wavelength range (400-1000 nm) and a high resolution of 1240$\times$1680$\times$200. We randomly selected and preprocessed 50 images, generating 276 patches of 256$\times$256$\times$70 for trainnig and 6 scenes for testing. Houston is a commonly-used HSI in remote sensing of size 349$\times$1905 $\times$144 in 380-1050 nm, which is used as one scene for cross-domain test of methods. The size of 7 testing scenes is 256$\times$2048$\times$70.
Real LASIS Dataset. We use absolute calibration data captured by the LASIS instrument (mentioned in Sec. 3.2) as our testing dataset. The dataset includes real interferograms captured in uniform light and corresponding ideal HSIs as the ground truth. The formats of all real data are shown in Tab. 1, with a size of $H$=120.
@@
width $W$=2048
channels $\Lambda=$70 $N$=221 $L$=256
center - - $l=$35
@@
$m_{2}$ $\mathbf{A},\bm{\beta},\bm{M}$ $\mathbf{G}$ 37.95 / 0.965
$m_{3}$ $\mathbf{A},\bm{\beta},\bm{M}$ $\mathbf{G},\mathbf{N}_{shot}$ 38.16 / 0.974
$m_{4}$ $\mathbf{A},\bm{\beta}$ 26.20 / 0.863
$m_{5}$ $\mathbf{A},\bm{\beta},\bm{M}$ 33.96 / 0.839
$m_{6}$ $\mathbf{A},\bm{\beta},\bm{M}$ $\mathbf{N}_{shot},\mathbf{N}_{read}$ 39.28 / 0.988
ours $\mathbf{A},\bm{\beta},\bm{M}$ $\mathbf{N}_{shot},\mathbf{N}_{read},e$ 39.46 / 0.992
Table 3: Validation of degradation model on real dataset. Among the models, $m_{2},m_{3}$ are equipped with Gaussian white noise [33] for comparison. Our model achieve the optimal performance.
@@ 4.2 Degradation Model Validation
To verify the effectiveness of our degradation model, we train the 5-stage IHRUT network with data simulated from our model along with 6 comparative models, and analyze the reconstruction performance on real testing dataset. Among these models, traditional model $m_{1}$ is used for guiding traditional reconstruction method instead of data synthesis. The models $m_{2}$ and $m_{3}$ include the introduction of Gaussian white noise model for comparison with our noise model, while $m_{4}$ to $m_{6}$ are ablation models of our model. The experimental results are shown in Tab. 4. Our degradation model significantly outperforms traditional models (+12 dB). Compared to noise models in $m_{2}$ and $m_{3}$, our sensor-wise noise model is more accurate (+1.3 dB). Furthermore, the results of $m_{4}$ to $m_{6}$ validate the effectiveness of each degradation component in the model, particularly the introduction of $e$ that enhances robustness (+0.18 dB). The comparative results validates the rationality and accuracy of our degradation model structure.
@@ 4.3 Reconstruction Experiments
Figure 4: Comparison of reconstructed HSI with error maps (10$\times$ amplified). We select Synthetic (a) Scene 2 from HSOD-BIT and (b) Scene 1 from Houston cropped by 256$\times$256, along with Real (c) Scene 3 from Calibration dataset cropped by 64$\times$64. The error map exhibits the spectral absolute error of IHI reconstruction results by 8 algorithms and IHRUT with 7 stages. The region within the box is chosen for analysis of the reconstructed spectra and zoom in for a more detailed examination.
@@
4.4 Ablation Study
On the basis of verifying the overall effectiveness of our method, this section focuses on the ablation analysis of the 5-stage IHRUT network. Here, we present a baseline network by removing all modules of our design in IHRUT. As shown in Tab. 4, during the progressive reintroduction of the designed modules for ablation studies, we find that the SPE mechanism plays a significant role in enhancing network performance, resulting in a gain over 5 dB, while momentum mechanism and cross-stage fusion serve supplementary functions. As shown in Fig. 5, the characteristics of SPEs in different subnets are not identical, while they all exhibit notably effective roles in the removal of stripe patterns. For further analysis, we replace SPE with channel attention and observe a notable decline (0.7dB) in performance, which validates the effectiveness of our SPE design.
@@ 4.4 Ablation Study
Figure 5: The result of incorporating SPE in the ablation exhibits the effectiveness of SPE in eliminating stripe patterns.

### v2

@@ Abstract
Interferometric Hyperspectral Imaging (IHI) is a critical technique for large-scale remote sensing tasks due to its advantages in flux and spectral resolution. However, IHI is susceptible to complex errors arising from imaging steps, and its quality is limited by existing signal processing-based reconstruction algorithms. Two key challenges hinder performance enhancement: 1) the lack of training datasets. 2) the difficulty in eliminating IHI-specific degradation components through learning-based methods. To address these challenges, we propose a novel IHI reconstruction pipeline. First, based on imaging physics and radiometric calibration data, we establish a simplified yet accurate IHI degradation model and a parameter estimation method. This model enables the synthesis of realistic IHI training datasets from hyperspectral images (HSIs), bridging the gap between IHI reconstruction and deep learning. Second, we design the Interferometric Hyperspectral Reconstruction Unfolding Transformer (IHRUT), which achieves effective spectral correction and detail restoration through a stripe-pattern enhancement mechanism and a spatial-spectral transformer architecture. Experimental results demonstrate the superior performance and generalization capability of our method. The code and are available at https://github.com/bit1120203554/IHRUT.
@@ 2.2 Degradation Modeling and Dataset Synthesis
Learning-based restoration [40, 55, 30] approaches typically draw their potential from data. However, the acquisition of large-scale real datasets is often challenging, and one solution is using degradation models as priors to generate synthetic datasets [24]. Common models includes Gaussian white noise and homoscedastic Gaussian noise [37, 1, 16], among others. Wei [41, 42] introduced a camera calibration-based, precise noise model that allows simulated data to yield low-light noise recovery akin to real-data training outcomes. Zhang [52, 53, 54] extended this to HSI denoising, presenting a realistic dataset and an HSI noise model. These methods provide inspirations for addressing the challenges in existing IHI calibration and generating training data.
@@ 2.3 Hyperspectral Image Reconstruction
Existing HSI reconstruction focus on tasks such as RGB or infrared-to-HSI conversion [18, 46], and compressed sensing imaging (SCI) [2, 23]. Deep learning approaches outperforms traditional model-driven methods [5, 47] in terms of performance and generalization. These methods are categorized into two types: end-to-end (E2E) and model-aided networks. E2E methods directly learn the reconstruction mapping , with examples including CNN-based TSA-Net [33] and HDNet [20], RNN-based BIRNAT [11], and transformer-based MST, MST++ , and CST [7, 8, 6]. Model-aided networks, exemplified by Deep Plug-and-Play (PnP) [56] and Deep Unfolding (DU), integrate iterative optimization with deep learning. Among these, DU adopts the same training strategy as E2E methods, thereby enhancing the network’s adaptability to iterations compared to PnP, and achieving superior restoration performance. Examples include GAP-Net [34], transformer-based DAUHST [9], Proximal Gradient Descent (PGD)-based RDLUF [13], and Memory-Augmented MAUN [19]. However, for IHI tasks, these algorithms have yet to be applied and optimized.
@@ 3.1 Degradation Modeling
The input light signal can be considered as a hyperspectral signal in a certain wavelength range $\lambda_{\text{min}},\lambda_{\text{max}}]$. Assuming expression of the signal in the form of wavenumber $\nu$ ($\nu=1/\lambda\in[1/\lambda_{\text{max}},1/\lambda_{\text{min}}]$) is $\mathbf{B}_{0}(\nu)\in\mathbb{R}^{H\times W\times N}$. According to theories of IHI [4], $\mathbf{B}_{0}$ is Fourier transform-related to the interferogram. However, due to the imaging system’s varying sensitivity to light in different frequencies, the actual captured spectral signal is the result of the original signal multiplied by a coefficient vector $\mathbf{A}$ in the spectral domain. Additionally, pose deviations in the interferometer lead to asymmetry in the interferogram, introducing a complex component as phase error into $\mathbf{A}$. We define $\mathbf{A}\in\mathbb{C}^{W\times N}$ as the absolute response coefficient, and the interferogram signal can be represented as:
@@ 4.1 Datasets and Settings
[…]
Real LASIS Dataset. We use absolute calibration data captured by the LASIS instrument (mentioned in Sec. 3.2) as our testing dataset. The dataset includes real interferograms captured in uniform light and corresponding ideal HSIs as the ground truth. The formats of all real data are shown in Tab. 1, with a size of HH=120.
@@
width WW=2048
channels $\Lambda=$70 NN=221 LL=256
center - - l=l=35
@@
$m_{2}$ $\mathbf{A},\bm{\beta}$ - 26.20 / 0.863
$m_{3}$ $\mathbf{A},\bm{\beta},\bm{M}$ - 33.96 / 0.839
$m_{4}$ $\mathbf{A},\bm{\beta},\bm{M}$ Gaussian Noise 37.95 / 0.965
$m_{5}$ $\mathbf{A},\bm{\beta},\bm{M}$ Standard Noise 38.82 / 0.987
$m_{6}$ $\mathbf{A},\bm{\beta},\bm{M}$ Our Noise (pixel-wise) 39.28 / 0.988
ours $\mathbf{A},\bm{\beta},\bm{M}$ Our Noise + $e$ 39.46 / 0.992
Table 3: Validation of degradation model on real dataset. Among the models, $m_{4}$ is equipped with Gaussian white noise [37] for, and $m_{5}$ with standard noise model ($\mathbf{N}_{shot},\mathbf{N}_{read}$) for comparison. Our model achieve the optimal performance.
@@ 4.2 Degradation Model Validation
To verify the effectiveness of our degradation model, we train the 5-stage IHRUT network with data simulated from our model along with 6 comparative models, and analyze the reconstruction performance on real testing dataset. Among these models, model $m_{1}$ is used for guiding traditional reconstruction. The models $m_{2}$ and $m_{3}$ serves as ablations of our optical parts (such as $\mathbf{M}$). $m_{4}$ and $m_{5}$ are equipped with Gaussian white noise and standard noise model ($\mathbf{N}_{shot},\mathbf{N}_{read}$) for comparison. With our design of pixel-wise noise modeling for the stripe-pattern noise distribution, $m_{6}$ outperforms standard noise model by 0.64dB. Furthermore, the introduction of $e$ effectively enhances robustness (+0.18 dB). The comparative results validates the rationality and accuracy of our degradation model structure.
@@ 4.3 Reconstruction Experiments
[…]
@@
4.4 Ablation Study
Under the validated overall efficacy of the proposed framework, ablation experiments are conducted in two stages. First, a comprehensive breakdown ablation is performed on every major component integrated in IHRUT. As reported in Table 4, the SPE mechanism emerges as the most critical contributor, delivering a performance gain exceeding 5 dB, whereas the momentum mechanism and cross-stage fusion play auxiliary roles in boosting quality of reconstruction.
Subsequently, we undertake a detailed comparative analysis of our SPE and Momentum Branch against channel-attention (Tab. 5 a,b,c) and existing momentum branches (Tab. 5 d,e). As illustrated in Tab. 5 and Fig. 5, the stripe-wise attention introduced by SPE yields superior flexibility, and the two successive weighting operations within our momentum pathway further elevate the overall performance.
Data Module a b c
$z_{k}-x_{k}$ $\text{CA}(\text{CA}\cdot\delta$+$m)$ $\text{CA}(\text{SPE}_{0}\cdot\delta$+$m)$ $\text{SPE}_{k}(\text{CA}\cdot\delta$+$m)$
PSNR/SSIM 38.55/0.984 38.66/0.984 38.60/0.985
Data Module d (like [3, 45]) e ours
$z_{k}-x_{k}$ $\text{CA}\cdot\delta$$\text{CA}(m)$ $\text{SPE}_{k}\cdot\delta$$\text{SPE}_{0}(m)$ $\text{SPE}_{k}(\text{SPE}_{0}\cdot\delta$+$m)$
PSNR/SSIM 38.31/0.983 38.40/0.985 39.46/0.992
Table 5: Analysis of the Data Modules with different components and momentum branches (e.g., $z_{k}-x_{k}$ in the table) in 5-stg IHRUT. Here, $\delta=\mathbf{y}-\mathbf{F}\mathbf{x}_{k}$ and $m=\mathbf{x}_{k}-\mathbf{x}_{k-1}$.
[…]
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Key R&D Program of China (2022YFC3300705), the National Natural Science Foundation of China (62331006, 62171038, and 62088101), and the Fundamental Research Funds for the Central Universities.